OpenStudy (anonymous):

In arithmetic sequence is given: \[\LARGE a_p=q \] \[\LARGE a_q=p \] Find \[\LARGE a_n=? \] \[\LARGE A)\quad p-q+n\] \[\LARGE B)\quad p-q-n\] \[\LARGE C)\quad-p+q+n\] \[\LARGE C)\quad p+q-n\] I can't even start it, leave alone to do it !!!

OpenStudy (anonymous):

Hmm. \[ a_{p} = a _{1}+(p-1)d=q\] \[ a_{q} = a _{1}+(q-1)d=p\] \[ a_{p} -a_{q}=(p-q)d=q-p\] from here d=-1 so:\[a _{n}=a _{1}-n+1\] But i can't find the conection with the options you have....

OpenStudy (anonymous):

@Kreshnik

OpenStudy (anonymous):

@myko Thank you very much. You Gave me a great help. ! \[\LARGE a_p=q -------[1]\] \[\LARGE a_q=p -------[2]\] \[\LARGE a_p=a_1+(p-1)d---[3] \] \[\LARGE a_q=a_1+(q-1)d----[4] \] Now from [1] and [3] if: \[\LARGE a_p=q \quad \quad and \quad \quad a_p=a_1+(p-1)d \quad \Longrightarrow \] \[\LARGE q=a_1+(p-1)d-----[5] \] And from [2] and [4] if: \[\LARGE a_q=p \quad \quad \quad and\quad \quad a_q=a_1+(q-1)d \quad \Longrightarrow\] \[\LARGE p=a_1+(q-1)d -----[6]\] You did well, from \[\LARGE a_p-a_q=[a_1+(p-1)d]-[a_1+(q-1)d]]= \] \[\LARGE =(p-1)d-(q-1)d=d[(p-1)-(q-1)]= \] \[\LARGE d(p-q) \Longrightarrow a_p-a_q=d(p-q) .\] And from [1] & [2] we know that: \[\LARGE a_p-a_q=q-p . \] \[ \LARGE \text{So:}\;\;\; d(p-q)=q-p \quad \Longleftrightarrow d=-1 \] Now we substitute d=-1 to the main formula: (This is what you've done, I just rewrote them longer :D) But before we substitute... from: [1] and [3] I did. \[\LARGE q=a_1+(p-1)d \quad \Longrightarrow a_1=q-(p-1)d \] since d=-1 ... \[\LARGE a_1=q-(p-1)d=q-(-p+1) \quad \Longrightarrow \] \[\LARGE a_1=q+p-1 \] And also from: [2] & [4] \[\LARGE p=a_1+(q-1)d \Longrightarrow a_1=p-(q-1)d \] \[ \LARGE \Longrightarrow a_1=p+q-1 \] so we get the same result... and now let's substitute \[\LARGE a_n=a_1+(n-1)d \Longrightarrow a_n=(p+q-1)+(1-n) \] \[\LARGE a_n=p+q-n \] I think this should be correct O_O ...

OpenStudy (anonymous):

can someone verify it .. if this is correct or not :O

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

you too ! :)

OpenStudy (amistre64):

an = p+q-n is correct

OpenStudy (anonymous):

thank you .